"Write the first four terms of the geometric sequence with the first term (a sub n = 16) and common ration r= -1/2.
Answer by venugopalramana(3286)

"Write the first four terms of the geometric sequence with the first term (a sub n = 16) and common ration r= -1/2.
A1=16
A2=A1*R=16*(-1/2)=-8
A3=A2*R=-8*(-1/2)=4
A4=A3*R=4*(-1/2)=-2